Angie is 25 years older than her daughter, Brooke. In 6 years Angie will be twice as old as Brooke. How old are they both now?
Answer by rothauserc(4718)

Let x be Brooke's age and y be Angie's age
:
1) y = x + 25
:
2) y + 6 = 2(x + 6)
:
substitute x + 25 for y in equation 2)
:
x + 25 + 6 = 2x + 12
:
x = 31 - 12 = 19
:
*******************************************
Brooke is 19 years old
Angie is 44 years old
:
Note that in 6 years Brooke is 25 years old
and Angie is 50 years old
********************************************
: